OpenStudy (anonymous):

solve and check extraeous solution 3square root 2x+4=12

OpenStudy (anonymous):

[3\sqrt{2x+4=12}\] divide by 3 \[\sqrt{2x+4} = 4\] multiply both sides 2 as exponent. it will become 2x +4 = 4\[^{2}\] 2x +4 = 16 2x = 16-4 2x = 12 divide by 2 x= 6

OpenStudy (anonymous):

can u do another?

OpenStudy (anonymous):

^ Ask away